# Rings in which one-sided strongly $\pi$-regular elements are strongly $\pi$-regular

### v1

@@ 3 Some classes of rings that are not right Dischinger and left-right asymmetry
The following result will be crucial for everything that follows. The last part of this result was first proved by Pace P. Nielsen and was communicated to the second author in 2012.
@@
In view of Theorem 3.2, it is easy to see $f\in Ra$. Suppose $f=f_{1}a^{k}$ such that $f_{1}\notin Ra$. Clearly $r_{R}(a^{k})\subseteq r_{R}(f)$. Let $h$ be a nonzero element in $R$ such that $fh=0$. This implies $f_{1}a^{k}h=0$. Since $f_{1}\notin Ra$, again by Theorem 3.2, $a^{k}h=0$ implies $h\in r_{R}(a^{k})$. ∎
Next, we identify all the nilpotent elements of the ring $R$ and also prove that they form a subring.
@@
In the notations of Lemma 2.12, $\rm{N}(R)=N^{r}_{R}(a)$. In particular, $R$ is an NR ring.
In view of Theorem 3.2, $N(R)\subseteq Ra$. It is also clear that $\mathbb{F}_{2}[a]\cap\rm{N}(R)=0$. Therefore, any nonzero nilpotent element of $R$ is of the form $r_{1}a$ and it must have at least one monomial in its support that is not in $\mathbb{F}_{2}[a]$. Then $ar_{1}$ is also nilpotent, and if $ar_{1}=0$, we are through. Note that any monomial of $r_{1}$ that was in $Rx$ will vanish from $ar_{1}$ either by cancellation or reduction. We continue this process of shifting $a$ to the left and in the process $ar_{k}\in\mathbb{F}_{2}[a]$ for some $k\in\mathbb{N}$. As $ar_{k}$ is nilpotent, so $ar_{k}=0$ implying that $\rm{N}(R)=N_{R}^{r}(a)$. ∎
@@
Suppose $h$ is a nonzero element in $R$ such that $hf=0$. As $f\neq 0$, by Corollary 3.3, we have $h=h_{1}a^{k}$ for some $k\in\mathbb{N}$ such that $h_{1}\notin Ra$. Since $h_{1}\notin Ra$ again by Corollary 3.3 , $r_{R}(h_{1})=0$. As $h_{1}a^{k}f=0$, so $a^{k}f=0$. If $n\in\mathbb{N}$ is the smallest integer such that $a^{n}f=0$, then it is clear that $l_{R}(f)=Ra^{n}$. ∎
Recall that an element $s$ of a ring $S$ is called a left zero-divisor if $r_{S}(s)\neq 0$.
@@
(i) The set of left zero divisors of $R$ is $Ra$.
(ii) The set of the right zero divisors of $R$ is $\cup_{n=1}^{\infty}r_{R}(a^{n})$.
(i) It is clear from Corollary 3.3 and from the fact that ra(1ax)=0ra(1-ax)=0 for every $r\in R$.
(ii) Direct consequence of Corollary 3.6
Since there is no infinite ascending chain of left ideals of $R$ of the form $Ra^{n}$, $n\in\mathbb{N}$, the following result is immediate from Corollary 3.6.
@@
A ring satisfying the ascending chain condition on left annihilators may not be right Dischinger.
@@
Azumaya in [2, Theorem 2] proved that a ring $S$ which has no infinite ascending chain of the form $l_{S}(s)\subseteq l_{S}(s^{2})\subseteq\ldots$, is left Dischinger. So $R$ is left Dischinger by Corollary 3.8. Lastly, suppose $Rf=Rf^{2}$ for some $f\in R$. As $R$ is left Dischinger, $fR=f^{2}R$ also implies that $f$ is strongly regular. So $f=eu$ for some unit $u$ and idempotent $e$ in $R$. But by [5, Theorem 2.3] only idempotents in $R$ are 0 and $1$, implying that either $f=0$ or $f$ is a unit. ∎
We will need the following result to prove that $R$ is a UU ring. This will also lead to a self-contained proof of the last part of Theorem 3.10.
@@
(i) If $f,\,g\in R$ are such that $fg=1$, then both $f$ and $g$ have $1$ in their supports.
(ii) No non-identity element of $\mathbb{F}_{2}[a]$ is invertible in $R$.
(iii) If $1+f$ is a unit in $R$, then $f\in Ra$.
(i) This is clear because the product of two monomials is $1$ iff both equal $1$.
(ii) Suppose to the contrary $1+f\in\mathbb{F}_{2}[a]$ is invertible in $R$ with inverse $1+g$, where $f\neq 0$ and $g$ don’t have $1$ in their supports. Also $g\in Ra$ by Lemma 3.11. Clearly, $g\notin\mathbb{F}_{2}[a]$ because otherwise $1+f\in\mathbb{F}_{2}[a]$ is invertible in $\mathbb{F}_{2}[a]$. This is not possible as $1$ is the only invertible element in the ring $\mathbb{F}_{2}[a]$. Now $(1+g)(1+f)=1$ implies $g+f=gf$. Let $M$ be the largest monomial in the support of $g$ that is not in $\mathbb{F}_{2}[a]$. Then $M$ is present in the support of $g+f$. We show that $M$ is not present in the support of $gf$. Suppose $M=ma^{k}$, where $m$ is a monomial in the support of $g$ and $a^{k}$, $k\in\mathbb{N}$, is a monomial in the support of $f$. Clearly $m\neq M$. So by the maximality of $M$, we have $m\prec M$. Thus $M=ma^{k}\prec m\prec M$, which is a contradiction.
(iii) Suppose $1+g$ is inverse of $1+f$. Then $f+g+fg=0$ and as $f$ and $g$ don’t have $1$ in their supports, so $f\in Ra$ by Lemma 3.11. ∎
Now, we prove another main result of this section.
@@
The proof follows on similar lines as that of Theorem 3.4. Let $1+f\in\rm{U}(R)$. By Lemma 3.12(iii), $f=f_{1}a$ for some $f_{1}\in R$. As $1+f=1+f_{1}a$ is a unit, so is $1+af_{1}$. Again by 3.12(iii), $af_{1}=f_{2}a$ for some $f_{2}\in R$. Similarly, $1+af_{2}$ is a unit. Proceeding on these lines $af_{k}\in\mathbb{F}_{2}[a]$, for some $k\in\mathbb{N}$, as in the proof of Theorem 3.4. As $1+af_{k}$ is a unit, so $af_{k}=0$ by Lemma 3.12(ii). This implies that $f\in N_{R}(a)$ and so $f$ is nilpotent. Thus $R$ is a UU ring. ∎
Part (i) of the following result was noted without proof in [5] also.
@@
(i) $\rm{J}(R)=0$.
(ii) $\rm{N}(R)+\rm{U}(R)\subseteq\rm{U}(R).$
(iii) For any $u\in\rm{U}(R)$, we have $u\rm{N}(R)\subseteq\rm{N}(R)$ and $\rm{N}(R)u\subseteq\rm{N}(R)$.
(i) For any nonzero $f\in R$ there exists $k\in\mathbb{N}$ such that $fx^{k}$ has a monomial in its support that is in $Rx$. So by Corollary 3.3, $fx^{k}$ is not nilpotent and by Theorem 3.13, $1+fx^{k}$ is not a unit. So $\rm{J}(R)=0$.
(ii) and (iii) follow from the facts that $\rm{N}(R)$ is a subring (Theorem 3.4) and units in $R$ are unipotents (Theorem 3.13). ∎
Camillo and Nielsen in [3] defined a ring $S$ to be right linearly McCoy if for any two linear polynomials $f,\,g\in S[x]$ with $fg=0$, there exists a nonzero $s\in S$ such that $fs=0$. A left linearly McCoy rings can be defined similarly. As a one-sided linearly McCoy ring is Dedekind-finite [3, Theorem 5.2], one may wonder if it is right Dischinger. This is not the case as we will prove that the ring $R$ is left and right linearly McCoy. We need the following result proved in [8, Proposition 3.3].
@@
Is a right or left McCoy ring right Dischinger?
Yao in [12] defined the class of weakly left duo rings as a generalization of left duo rings. In Corollary 2.8, we prove that a weakly left duo ring with identity is right Dischinger. But we don’t know the answer to the following question.
@@
Is a weakly left duo ring left Dischinger?
@@
Is a left quasi-duo ring a left Dischinger?
In Corollary 3.8, we proved that $R$ satisfies the ascending chain condition on left annihilators. In view of Corollary 2.14, $_{R}R$ does not have a finite uniform dimension. This can also be seen directly in view of the following result.
@@
Both 1ax1-ax and $a$ are nonzero elements of $R$, $R(1-ax)\cap Ra=0$ and $l_{R}(a)=0$. So the result follows from Lemma 3.21. ∎
The following question is natural to ask.
@@ 4 Acknowledgements
We are thankful to Pace P. Nielsen for the interesting questions he raised in private communication.

### v2

@@ 3 Some classes of rings that are not right Dischinger and left-right asymmetry
The following result will be crucial for everything that follows. The last part of this result was first proved by Nielsen and was communicated to the second author in 2012.
@@
In view of Theorem 3.2, it is easy to see $f\in Ra$. Suppose $f=f_{1}a^{k}$ such that $f_{1}\notin Ra$. Clearly $r_{R}(a^{k})\subseteq r_{R}(f)$. Let $h$ be a nonzero element in $R$ such that $fh=0$. This implies $f_{1}a^{k}h=0$. Since $f_{1}\notin Ra$, again by Theorem 3.2, $a^{k}h=0$ implies $h\in r_{R}(a^{k})$. ∎ Next, we identify all the nilpotent elements of the ring $R$ and also prove that they form a subring.
@@
In the notations of Lemma 2.12, $\rm{N}$$(R)$ = $\rm{N}$$^{r}_{R}$$(a)$. In particular, $R$ is an NR ring.
In view of Theorem 3.2, $\rm{N}$$(R)\subseteq Ra$. It is also clear that $\mathbb{F}_{2}[a]\cap\rm{N}$$(R)=0$. Therefore, any nonzero nilpotent element of $R$ is of the form $r_{1}a$ and it must have at least one monomial in its support that is not in $\mathbb{F}_{2}[a]$. Then $ar_{1}$ is also nilpotent, and if $ar_{1}=0$, we are through. Note that any monomial of $r_{1}$ that was in $Rx$ will vanish from $ar_{1}$ either by cancellation or reduction. We continue this process of shifting $a$ to the left and in the process $ar_{k}\in\mathbb{F}_{2}[a]$ for some $k\in\mathbb{N}$. As $ar_{k}$ is nilpotent, so $ar_{k}=0$ implying that $\rm{N}$$(R)=\rm{N}$$^{r}_{R}$$(a)$
@@
Suppose $h$ is a nonzero element in $R$ such that $hf=0$. As $f\neq 0$, by Corollary 3.3, we have $h=h_{1}a^{k}$ for some $k\in\mathbb{N}$ such that $h_{1}\notin Ra$. Since $h_{1}\notin Ra$ again by Corollary 3.3 , $r_{R}(h_{1})=0$. As $h_{1}a^{k}f=0$, so $a^{k}f=0$. If $n\in\mathbb{N}$ is the smallest integer such that $a^{n}f=0$, then it is clear that $l_{R}(f)=Ra^{n}$. ∎ Recall that an element $s$ in a ring $S$ is called a left zero-divisor if $r_{S}(s)\neq 0$.
@@
(1) The set of left zero divisors of $R$ is $Ra$.
(2) The set of the right zero divisors of $R$ is $\cup_{n=1}^{\infty}r_{R}(a^{n})$.
(1) It is clear from Corollary 3.3 and from the fact that ra(1ax)=0ra(1-ax)=0 for every $r\in R$.
(2) This follows from Corollary 3.6. ∎ Since there is no infinite ascending chain of left ideals of $R$ of the form $Ra^{n}$, $n\in\mathbb{N}$, the following result is immediate from Corollary 3.6.
@@
A ring satisfying the ascending chain condition on left annihilators of subsets may not be right Dischinger.
@@
Azumaya in [2, Theorem 2] proved that a ring $S$ which has no infinite ascending chain of the form $l_{S}(s)\subseteq l_{S}(s^{2})\subseteq\ldots$, is left Dischinger. So $R$ is left Dischinger by Corollary 3.8. Lastly, suppose $Rf=Rf^{2}$ for some $f\in R$. As $R$ is left Dischinger, $fR=f^{2}R$ also implies that $f$ is strongly regular. So $f=eu$ for some unit $u$ and idempotent $e$ in $R$. But by [5, Theorem 2.3] only idempotents in $R$ are 0 and $1$, implying that either $f=0$ or $f$ is a unit. ∎ We will need the following result to prove that $R$ is a UU ring. This will also lead to a self-contained proof of the last part of Theorem 3.10.
@@
(1) If $f,\,g\in R$ are such that $fg=1$, then both $f$ and $g$ have $1$ in their supports.
(2) No non-identity element of $\mathbb{F}_{2}[a]$ is invertible in $R$.
(3) If $1+f$ is a unit in $R$, then $f\in Ra$.
(1) This is clear because the product of two monomials is $1$ iff both equal $1$.
(2) Suppose to the contrary $1+f\in\mathbb{F}_{2}[a]$ is invertible in $R$ with inverse $1+g$, where $f\neq 0$ and $g$ don’t have $1$ in their supports. Also $g\in Ra$ by Lemma 3.11. Clearly, $g\notin\mathbb{F}_{2}[a]$ because otherwise $1+f\in\mathbb{F}_{2}[a]$ is invertible in $\mathbb{F}_{2}[a]$. This is not possible as $1$ is the only invertible element in the ring $\mathbb{F}_{2}[a]$. Now $(1+g)(1+f)=1$ implies $g+f=gf$. Let $M$ be the largest monomial in the support of $g$ that is not in $\mathbb{F}_{2}[a]$. Then $M$ is present in the support of $g+f$. We show that $M$ is not present in the support of $gf$. Suppose $M=ma^{k}$, where $m$ is a monomial in the support of $g$ and $a^{k}$, $k\in\mathbb{N}$, is a monomial in the support of $f$. Clearly $m\neq M$. So by the maximality of $M$, we have $m\prec M$. Thus $M=ma^{k}\prec m\prec M$, which is a contradiction.
(3) Suppose $1+g$ is inverse of $1+f$. Then $f+g+fg=0$ and as $f$ and $g$ don’t have $1$ in their supports, so $f\in Ra$ by Lemma 3.11. ∎ Now, we prove another main result of this section.
@@
The proof follows on similar lines as that of Theorem 3.4. Let $1+f\in\rm{U}$$(R)$. By Lemma 3.12(3), $f=f_{1}a$ for some $f_{1}\in R$. As $1+f=1+f_{1}a$ is a unit, so is $1+af_{1}$. Again by 3.12(3), $af_{1}=f_{2}a$ for some $f_{2}\in R$. Similarly, $1+af_{2}$ is a unit. Proceeding on these lines $af_{k}\in\mathbb{F}_{2}[a]$, for some $k\in\mathbb{N}$, as in the proof of Theorem 3.4. As $1+af_{k}$ is a unit, so $af_{k}=0$ by Lemma 3.12(2). This implies that $f\in N_{R}(a)$ and so $f$ is nilpotent. Thus $R$ is a UU ring. ∎
Part (1) of the following result was noted without proof in [5] also.
@@
(1) $\rm{J}$$(R)=0$.
(2) $\rm{N}$$(R)+\rm{U}$$\subseteq\rm{U}(R).(R)$.
(3) For any $u\in\rm{U}$$(R)$, we have $u\rm{N}$$\subseteq\rm{N}(R)(R)$ and $\rm{N}$$u\subseteq\rm{N}(R)(R)$.
(1) For any nonzero $f\in R$ there exists $k\in\mathbb{N}$ such that $fx^{k}$ has a monomial in its support that is in $Rx$. So by Corollary 3.3, $fx^{k}$ is not nilpotent and by Theorem 3.13, $1+fx^{k}$ is not a unit. So $\rm{J}$$(R)=0$.
(2) and (3) follow from the facts that $\rm{N}$$(R)$ is a subring (Theorem 3.4) and units in $R$ are unipotents (Theorem 3.13). ∎ Camillo and Nielsen in [3] defined a ring $S$ to be right linearly McCoy if for any two linear polynomials $f,\,g\in S[x]$ with $fg=0$, there exists a nonzero $s\in S$ such that $fs=0$. A left linearly McCoy rings can be defined similarly. As a one-sided linearly McCoy ring is Dedekind-finite [3, Theorem 5.2], one may wonder if it is right Dischinger. This is not the case as we will prove that the ring $R$ is left and right linearly McCoy. We need the following result proved in [8, Proposition 3.3].
@@
Is a right or left McCoy ring with identity right Dischinger?
Yao in [12] defined the class of weakly left duo rings as a generalization of left duo rings. In Corollary 2.8, we proved that a weakly left duo ring with identity is right Dischinger. But we don’t know the answer to the following question.
@@
Is a weakly left duo ring with identity left Dischinger?
@@
Is a left quasi-duo ring with identity a left Dischinger?
In Corollary 3.8, we proved that $R$ satisfies the ascending chain condition on left annihilators of subsets of $R$. In view of Corollary 2.14, $_{R}R$ does not have a finite uniform dimension. This can also be seen directly in view of the following result.
@@
Both 1ax1-ax and $a$ are nonzero elements of $R$, $R(1-ax)\cap Ra=0$ and $l_{R}(a)=0$. So the result follows from Lemma 3.21. ∎ The following question is natural to ask.
@@ 4 Acknowledgements
We are thankful to Pace P. Nielsen for many interesting questions that he raised in private communication. We are also grateful to him for reading the paper and for many useful comments that improved the quality of the paper.